OpenStudy (anonymous):

(12[ ]6)[ ](4[ ]2)=9 in the bracket fill in with the math symbol +,-,x,/ only use the symbol once

OpenStudy (anonymous):

multiply, divide, multiply so... 12*6 = 72 72/ (4*2) 4*2 = 8 72/8 = 9

OpenStudy (anonymous):

you can only use the symbol once

OpenStudy (anonymous):

(12+6)/(4-2)

OpenStudy (anonymous):

ok ty